OpenStudy (anonymous):

Dividing Radical Expressions Rationalize the denominator of each expression.

OpenStudy (anonymous):

\[\frac{ \sqrt[4]{2x} }{ \sqrt[4]{3x^2} }\]

OpenStudy (anonymous):

@sammixboo @DanJS @wio Can u help me plzz?

OpenStudy (anonymous):

@One098 @myininaya @zepdrix

OpenStudy (ajspeller):